A balance and torque experiment is conducted with an UN-balanced ruler.

Length of the ruler: 1 m

Mass of the ruler: 193 g

The pivot is at the 35-cm mark.

A 46-g mass is located at the 82-cm mark of the ruler.

We want to balance the ruler by placing another mass at the 19-cm mark.

What should be magnitude of this additional mass?
Answer #1

Balancing the torque about the pivot point,

M x (35 - 19) = 193(50 - 35) + 46(82 - 35)

M x 16 = 2895 + 2162

M = 316.06 g

Comment in case any doubt please rate my answer....